a) Arrange the following species in terms of increasing CL-O bond energy ( lowest to highest, use "=" if you believe two have the same energy).

CLOH, CLO4-, CLO3-, CLO2-

b) What is the average charge (formal) on the oxygen in each compound? (Show formula for Average formal charge)

c) Which compound do you think is more basic? why?
Answer #1

Answer :

Leon - и око ce - oli o no eo - onl=0 urce=o oceno como ö :0=1=0 & ö= o=g=0 0 = le=0

Molecule resonance forms no. of total ClO bond in structure Cl-O bond order = total no. of ClO bond / resonance structure bond energy rank
HOCl 1 1 1/1 = 1 4 (lowest )
ClO​​​​2​​​- 2 3 3/2 = 1.5 3
ClO​​​​3 - 3 5 5/3 = 1.66 2
ClO​4 - 4 7 7/4 = 1.75 1(highest)

Since bond order is directly proportional to bond energy ( higher is the bond order higher is the energy required to break that bond ) .

b) : average charge on oxygen in each compound = charge / no. of oxygen atom

Compound charge on compound no. of oxygen atom average charge on each oxygen
HOCl 0 1 0/1 = 0
ClO​​​​2 - -1 2 -1/2 = -0.5
ClO​​​​3 - -1 3 -1/ 3 = -0.33
ClO​​​​4 - -1 4

-1/4 = -0.25

C) : more basic is HOCl .

For acting as a Lewis base Cl has to donate the electrons . Maximum basic character is observed when Cl has maximum number of lone pair of electrons .

Also , as O is electronegative atom than Cl , it will show -I effect and will decrease the lone pair donation capacity of Cl . So, maximum will be the no. of O atom with Cl , lesser will be it's basicity .

According to both reason basicity order is :

HOCl > ClO​​​​2 - > ClO​​​​3 - > ClO​​​​4 - .